OpenStudy (anonymous):

How do you solve this? y=2x²-8

OpenStudy (anonymous):

what are you solving for?

OpenStudy (anonymous):

x or y?

OpenStudy (anonymous):

Well I need to find the intercepts and use them to graph the equation.

OpenStudy (anonymous):

Perhaps you are asking where y = 0? Which would be:\[y = 0 = 2x ^{2} - 8\] If that is what you are looking for, then first factor out the "2" and then factor x^2 - 4. You will get y = 0 = 2(x - x1)(x + x1), but it won't be the equation from @Dualknight98 as he is a bit off on the x1. See if you can attempt getting the x1 value. Tutors, please allow the student to try the problem.

OpenStudy (anonymous):

Hint on the factoring: find 2 numbers (positive and negative the SAME number) that when multiplied will equal negative 4. An even easier way to look at it is what is the positive square root of 4? That will be your "+" x1 and your "-" x1 for my equation above.

OpenStudy (anonymous):

I know I need to find two numbers inside the parenthesis that equal -8 when multiplied... but at the same time eliminate them selves in the process.. that's where I'm stuck

OpenStudy (anonymous):

No, you don't need to do what you just said. You you need to work with -4, not -8. Re-read my posts.

OpenStudy (anonymous):

Oooh I think I got it. I came up with 2(x-2)(x+2)

OpenStudy (anonymous):

Very excellent! Now, we can go to the next step of intercepts!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First, as you might already know, this is a parabola opening upward, so the graph will look like a bowl where the sides go up forever. There will be 2 places where the graph crosses the x-axis (which is where y =0) which are the x-intercepts. The y-intercept is where x=0 and is the minimum of the "bowl". For x < -2, and for x > 2, the y values are positive. For -2 < x < 2, the y values are negative.

OpenStudy (anonymous):

So, points (-2, 0) and (2, 0) are the x-intercepts. Can you tell me what the y-intercept is?

OpenStudy (anonymous):

y=-8

OpenStudy (anonymous):

Great! Looks like you are catching on very quickly! I found a great website that will graph if you would like to see it. I plugged your equation in and it shows your graph. Would you like to see it?

OpenStudy (anonymous):

yes please. thanks!

OpenStudy (anonymous):

Looks like:

OpenStudy (anonymous):

That "bottom" is at -8, but the scal is a little weird. This is from the website in the corner. Pretty easy to use. I had to put the equation in as: y = 2 * x^2 - 8 with that asterisk.

OpenStudy (anonymous):

That website is http://www.meta-calculator.com/online/ So, all good, now?

OpenStudy (anonymous):

thank you for all your help!

OpenStudy (anonymous):

Nice working with you! If you like my help, then hit "best response".

OpenStudy (anonymous):

You're welcome :]